OpenStudy (anonymous):

simplify 2^3√162 please explain how i should do this!! ;)

OpenStudy (ajprincess):

vat is 2^3?

OpenStudy (anonymous):

like its a 2 then an exponent 3 then the root thingie then 162

OpenStudy (ajprincess):

I mean vat is 2^3=?

OpenStudy (ajprincess):

the value of it.

OpenStudy (anonymous):

uhhhh...... 8?

OpenStudy (ajprincess):

ok nw √162 =?

OpenStudy (anonymous):

its not even... its a decimal... ;(

OpenStudy (anonymous):

*cough* factor *cough*

OpenStudy (ajprincess):

√162=√2*9*9 =9√2

OpenStudy (ajprincess):

i meant this.

OpenStudy (anonymous):

thats not the answer... ;(

OpenStudy (ajprincess):

 yep I knw. multiply 8 by 9√2 to get the answer.